Project

General

Profile

05 15 14

Attending: Eric, Nate, Craig Pavan, Michael, Kanika, Satish, Dominick, Jan, and Chris.

Craig is working on the FTS request for SCD to optimize our FTS resources.
Will send something around by the end of the week so that you can check for things I’m missing.

Need to update this page:
https://cdcvs.fnal.gov/redmine/projects/nova_sam/wiki/Official_Datasets

Code status: Not much to report since Monday. Started building the newest Art…
—> Backported preselection change on Tuesday.

Disk status: Some old cosmic files have been removed (7.1 TB)
—> On the Wiki: https://cdcvs.fnal.gov/redmine/projects/novaart/wiki/Archiving_data_and_access_to_data_on_tapes

Weird FTS behavior..
—> Not picking up certain data_tier files.
—> Nate has an idea of what is going on.. Eric and Nate bumped heads — duplicates causing confusion?

Production status:

— 4DB MC  (tight preselection)
—> Done with . (about 1/2 of the PID part missing)
—> Announced that about 1/2 ready

loose preselection status:

—>  Cosmic Data 
—> Ran all PID
—> Ran a lot of LEMSUM
—> LEM taking much longer than they used to (24 hours compared to 2 hours)
—> Need the number of slices that pass loose preselection compared to tight preselection.
—> CAF’s within a few hours. Hopefully CAF’s tomorrow morning.
—>  Cosmic 4DB MC 
—> None of the CRY files have been processed yet.
—> Need to be resubmitted.
—> Nate and Eric will coordinate and get it done today.
—>  GENIE MC     
—> 1/2 PID processed ( done off-site)
—> Eric checking on log files now.
—>

ND w/ overlays:

—>  20 jobs with 1000 events each (final testing)  
—> Only DAQ info in the overlaid files
—> Ran on Fermigrid.
—> Metadata status? Ready to go…
—> Offsite is okay now – solved post script problem.
—> Rock files fetched from SAM — ifdh_fetch
—> Feature upgrade will make it possible to store files at the local site.